# Asymptotic behaviors and dynamics of degenerate and mixed solitons for the coupled Hirota system with strong coherent coupling effects

### v1

@@ Abstract
In this work, we study the asymptotic behaviors and dynamics of degenerate and mixed solitons for the coupled Hirota system with strong coherent coupling effects in the isotropic nonlinear medium. Using the binary Darboux transformation, we derive the solutions to represent the degenerate solitons with two eigenvalues that are conjugate to each other. We obtain three types of degenerate solitons and provide their asymptotic expressions. Notably, these degenerate solitons exhibit time-dependent velocities, and the relative distance between the two asymptotic solitons increases logarithmically with the higher-order perturbation parameter $|\varepsilon|$ increasing. We also asymptotically reveal four interaction mechanisms between a degenerate soliton and a bell-shaped soliton: (1) elastic interaction with a position shift; (2) inelastic interaction for the degenerate soliton but elastic for the bell-shaped one; (3) elastic interaction for the degenerate soliton but inelastic for the bell-shaped one; and (4) the coherent interaction during a longer interaction region and elastic interaction based on specific parameter conditions. Besides, we analyze a special degenerate vector soliton that exhibits significant coherence effects, and numerically study the relationship between the robustness of such solitons and parameter $\varepsilon$. Our results indicate that $\varepsilon$ significantly affects the coherence of these solitons, and their robustness decreases when $|\varepsilon|$ increases.
@@ 1. Introduction
Solitons, which preserve their shapes during the propagation [1, 2], are extensively studied in nonlinear optics [3, 4, 5], Bose-Einstein condensates [6], plasma physics [7], and molecular biology [8]. The focusing nonlinear Schrödinger (NLS) equation governs the propagation of optical solitons in the picosecond regime, balancing group velocity dispersion with self-phase modulation [9]. Experimental observations and theoretical predictions of optical soliton propagation in optical fibers [1, 3, 4] have strongly driven the development of both mathematical and physical models for solitons. The focusing NLS equation is known to admit $N$-soliton solutions, which describe the elastic interactions of $N$ solitons in an ideal optical Kerr medium, where the reflection coefficient has $N$ simple poles according to the inverse scattering transform [10, 11]. Namely, an $N$-soliton solution corresponds to $N$ distinct eigenvalues of the linear spectral problem. As the time evolution $t\to\pm\infty$, an $N$-soliton solution can be viewed as a superposition of $N$ individual solitons, which separate linearly with $t$ and exhibit no interaction force [10, 11].
When the eigenvalues or their conjugates converge to the same value, degenerate soliton solutions, i.e., multi-pole solutions in the context of the inverse scattering transform, can be obtained [11, 12, 13, 14]. Distinct from the usual $N$-soliton interactions, degenerate solitons of the NLS equation can display intense interactions in the near-field region, where the relative distance between interacting solitons increases logarithmically with $|t|$ [13, 14]. In addition, all the asymptotic soliton branches of a degenerate soliton have the equal amplitudes and group velocities [13, 14]. In practical optical communication systems, researchers strive to create arrays of identical fundamental solitons. However, even small disturbances can cause interactions among neighboring solitons. Degenerate soliton solutions serve as the useful models for describing the behavior of multiple chirped pulses with the identical amplitudes and group velocities, especially when they are introduced without any phase difference [15, 16, 17]. In recent years, degenerate soliton solutions have been investigated in coupled NLS systems, revealing some fascinating dynamical properties [17, 18, 19].
@@
Coupled NLS systems are highly significant in both theory and practice for describing the behavior of multiple modes, frequencies, and polarizations in optical fibers and related structures [20, 21, 22]. These systems enable intensities transfer among additional degrees of freedom and generate a wide variety of vector solutions [20, 21, 22]. To describe the propagation of orthogonally polarized optical waves in an isotropic medium, the coupled NLS system with negative coherent coupling has been studied, which exhibits abundant coherent nonlinear phenomena [23, 24]. Additionally, for higher pulse input powers, it is important to incorporate higher-order effects into the basic NLS systems [9, 25, 26, 27]. In nonlinear optics, the third-order dispersion and self-steepening effects are used to describe the propagation of ultrashort pulses in optical fibers [9, 25, 26, 27]. Consequently, researchers have introduced the coupled Hirota system with higher-order effects [28, 29, 30]
${\rm i}q_{1,t}+q_{1,xx}+2\left(|q_{1}|^{2}+2|q_{2}|^{2}\right)q_{1}-2q_{1}^{*}q_{2}^{2}+{\rm i}\varepsilon q_{1,xxx}$
$+6{\rm i}\varepsilon\left(|q_{1}|^{2}+|q_{2}|^{2}\right)q_{1,x}+6{\rm i}\varepsilon\left(q_{1}q_{2}^{*}-q_{2}q_{1}^{*}\right)q_{2,x}=0,$ (1a)
${\rm i}q_{2,t}+q_{2,xx}+2\left(|q_{2}|^{2}+2|q_{1}|^{2}\right)q_{2}-2q_{2}^{*}q_{1}^{2}+{\rm i}\varepsilon q_{2,xxx}$
$+6{\rm i}\varepsilon\left(|q_{1}|^{2}+|q_{2}|^{2}\right)q_{2,x}+6{\rm i}\varepsilon\left(q_{2}q_{1}^{*}-q_{1}q_{2}^{*}\right)q_{1,x}=0,$ (1b)
@@
where $q_{1}=q_{1}(x,t)$ and $q_{2}=q_{2}(x,t)$ represent the slowly varying complex amplitudes of two interacting optical modes, the functions depend on the propagation variable $x$ and the time variable $t$, and the asterisk denotes the complex conjugate. The terms $|q_{1}|^{2}q_{1}$ and $|q_{2}|^{2}q_{2}$ account for self-phase modulation, while the terms $|q_{1}|^{2}q_{2}$ and $|q_{2}|^{2}q_{1}$ describe cross-phase modulation, and the terms $q_{1}^{*}q_{2}^{2}$ and $q_{2}^{*}q_{1}^{2}$ represent coherent coupling effects governing the energy exchange between the two modes of the fiber, and $\varepsilon$ is a higher-order perturbation parameter. When $\varepsilon\neq 0$, System (1) incorporates high-order effects such as third-order dispersion and self-steepening, which facilitate the description of orthogonally polarized optical pulse propagation in an isotropic medium. System (1) can be symplified into the matrix form of
${\rm i}{\bf Q}_{t}+{\bf Q}_{xx}+2{\bf QQ}^{\dagger}{\bf Q}+{\rm i}\varepsilon{\bf Q}_{xxx}+6{\rm i}\varepsilon{\bf QQ}^{\dagger}{\bf Q}_{x}={\bf 0},$ (2)
@@
where ${\bf Q}=\begin{pmatrix}q_{1}&q_{2}\\ -q_{2}&q_{1}\end{pmatrix}$ and “ $\dagger$ ” denotes the conjugate transpose, which is also called the matrix Hirota equation. It is noted that System (1) is unchanged through $x\rightarrow-x$ and $\varepsilon\rightarrow-\varepsilon$. Namely, we only need to consider the absolute value of higher-order perturbation parameter $\varepsilon$. Therefore, we always set $\varepsilon>0$ in this paper.
@@
For System (1), Ref. [28] has explored one- and two-hump solitons, bright and dark rogue waves, and bright and dark breathers with respect to a single spectral parameter using the Darboux transformation (DT) and generalized DT. Ref. [29] has studied the interactions of vector breathers with two spectral parameters, while Ref. [30] has reported the hybrid structures of rogue waves and breathers by the generalized DT.
@@
The aim of this paper is to investigate the asymptotic behaviors and coherent dynamics for the higher-order effects of degenerate and mixed solitons [^1] Mixed solitons describe interactions between degenerate solitons and other bell-shaped solitons. for System (1). The structure of this paper is outlined as follows: In Section 2, we will construct the $N$th-order binary DT and derive the determinant representation of soliton solutions for System (1). Section 3 will focus on deriving three types of degenerate solitons and their asymptotic expressions based on the binary DT. In Section 4, we will analyze the asymptotic expressions of mixed solitons before and after interactions, describing the four interaction mechanisms. In Section 5, we will discuss the interaction properties of mixed solitons based on both asymptotic and graphical analyses. In Section 6, we will analyze a special degenerate vector soliton with significant coherence effects and numerically investigate the relationship between the soliton’s robustness and parameter $\varepsilon$. Section 7 will conclude the paper.
@@ 2. The $N$th-order binary DT
Based on the Ablowitz-Kaup-Newell-Segur inverse scattering formulation [31], System (1) is completely integrable and admit the $4\times 4$ Lax pair [28, 29, 30]
$\Psi_{x}=U(\lambda;W)\Psi,\qquad\Psi_{t}=V(\lambda;W)\Psi,$ (3)
where $\Psi$ is a $4\times 1$ vector complex differentiable eigenfunction of $x$, $t$ and the complex spectral parameter $\lambda$, and $U$ and $V$ have the following forms of
$U(\lambda;W)={\rm i}\lambda\Omega+W,$
$V(\lambda;W)=4{\rm i}\varepsilon\Omega\lambda^{3}+2\left(2\varepsilon W+{\rm i}\Omega\right)\lambda^{2}+2\left({\rm i}\varepsilon W^{2}\Omega+W+{\rm i}\varepsilon W_{x}\Omega\right)\lambda$
$+2\varepsilon W^{3}+{\rm i}W^{2}\Omega-\varepsilon W_{xx}+{\rm i}W_{x}\Omega+\varepsilon\left(W_{x}W-WW_{x}\right),$
@@
$\Omega=\begin{pmatrix}-{\bf I}_{2\times 2}&{\bf 0}_{2\times 2}\\ {\bf 0}_{2\times 2}&{\bf I}_{2\times 2}\end{pmatrix},\qquad W=\begin{pmatrix}{\bf 0}_{2\times 2}&{\bf Q}\\ -{\bf Q}^{{\dagger}}&{\bf 0}_{2\times 2}\end{pmatrix},$
${\bf I}_{2\times 2}$ and ${\bf 0}_{2\times 2}$ as the $2\times 2$ identity matrix and zero matrix, respectively. From Lax Pair (3), we find that the compatibility condition $U_{t}-V_{x}+UV-VU=0$ is equivalent to System (1).
@@
with the superscript “ $-1$ ” denoting the inverse of a matrix. Hence, supposing that $\Psi_{j}=\left(\psi_{j,1},\psi_{j,2},\psi_{j,3},\psi_{j,4}\right)^{T}$ represents a vector eigenfunction of Lax Pair (3) corresponding to the complex eigenvalue $\lambda_{j}$ $(j=1,2,3,\cdots,N)$, it follows that $S\Psi_{j}=(-\psi_{j,2},\psi_{j,1},-\psi_{j,4},\psi_{j,3})^{T}$ is also a vector eigenfunction of Lax Pair (3) at $\lambda=\lambda_{j}$, where $N$ is a positive integer, $\psi_{j,k}$’s $(k=1,2,3,4)$ are the complex functions with respect to $x$ and $t$, and the superscript “ $T$ ” denotes the transpose operation on a vector/matrix. Setting that $\mathcal{H}_{j}=\left(\begin{array}[]{cc}\psi_{j,1}&-\psi_{j,2}\\ \psi_{j,2}&\psi_{j,1}\end{array}\right)$, $\mathcal{Y}_{j}=\left(\begin{array}[]{cc}\psi_{j,3}&-\psi_{j,4}\\ \psi_{j,4}&\psi_{j,3}\end{array}\right)$ and $\Gamma_{j}=\left(\Psi_{j},S\Psi_{j}\right)=\left(\begin{array}[]{c}\mathcal{H}_{j}\\ \mathcal{Y}_{j}\end{array}\right)$, we deduce that $\Gamma_{j}$ is a $4\times 2$ matrix eigenfunction of Lax Pair (3) at $\lambda=\lambda_{j}$, i.e.,
$\Gamma_{jx}=U(\lambda_{j};W)\Gamma_{j},\qquad\Gamma_{jt}=V(\lambda_{j};W)\Gamma_{j}.$ (5)
@@
where ${\bf Q}[N]=\left(\begin{array}[]{cc}q_{1}[N]&q_{2}[N]\\ -q_{2}[N]&q_{1}[N]\end{array}\right)$, $[N]$ indicates the $N$th iteration with respect to $q_{1}$ and $q_{2}$, $\varUpsilon_{1}[N]$ and $\varUpsilon_{2}[N]$ are both $2\times 2N$ matrices, which can be respectively represented as
$\varUpsilon_{1}[N]=\left(\begin{array}[]{ccccccc}\psi_{11}&-\psi_{12}&\psi_{21}&-\psi_{22}&\cdots&\psi_{N,1}&-\psi_{N,2}\\ \psi_{12}&\psi_{11}&\psi_{22}&\psi_{21}&\cdots&\psi_{N,2}&\psi_{N,1}\end{array}\right)=\left(\mathcal{H}_{1},\mathcal{H}_{2},\cdots,\mathcal{H}_{N}\right),$
$\varUpsilon_{2}[N]=\left(\begin{array}[]{ccccccc}\psi_{13}&-\psi_{14}&\psi_{23}&-\psi_{24}&\cdots&\psi_{N,3}&-\psi_{N,4}\\ \psi_{14}&\psi_{13}&\psi_{24}&\psi_{23}&\cdots&\psi_{N,4}&\psi_{N,3}\end{array}\right)=\left(\mathcal{Y}_{1},\mathcal{Y}_{2},\cdots,\mathcal{Y}_{N}\right),$
@@
$\varOmega[N]=\left(\begin{array}[]{cccc}\varOmega\left(\Gamma_{1},\Gamma_{1}\right)&\varOmega\left(\Gamma_{1},\Gamma_{2}\right)&\cdots&\varOmega\left(\Gamma_{1},\Gamma_{N}\right)\\ \varOmega\left(\Gamma_{2},\Gamma_{1}\right)&\varOmega\left(\Gamma_{2},\Gamma_{2}\right)&\cdots&\varOmega\left(\Gamma_{2},\Gamma_{N}\right)\\ \vdots&\vdots&\ddots&\vdots\\ \varOmega\left(\Gamma_{N},\Gamma_{1}\right)&\varOmega\left(\Gamma_{N},\Gamma_{2}\right)&\cdots&\varOmega\left(\Gamma_{N},\Gamma_{N}\right)\\ \end{array}\right),$ (13)

### v2

@@ Abstract
In this work, we investigate the asymptotic behaviors and dynamics of degenerate and mixed solitons in the coupled Hirota system with strong coherent coupling effect. Through the binary Darboux transformation, we obtain three types of degenerate solitons and their asymptotic expressions. These degenerate solitons admit time-dependent velocities, and the relative distance between the two asymptotic solitons logarithmically increases with the increase of the higher-order perturbation parameter $|\varepsilon|$. We also reveal four mechanisms of interaction between degenerate solitons and bell-shaped solitons: (1) elastic interaction with a position shift; (2) inelastic interaction for the degenerate soliton but elastic for the bell-shaped one; (3) elastic interaction for the degenerate soliton but inelastic for the bell-shaped one; and (4) elastic interaction based on coherent interaction under specific parameter conditions. Furthermore, we analyze a special degenerate vector soliton with strong coherent coupling effects, and through numerical studies, we investigate the relationship between the soliton’s robustness and the parameter $\varepsilon$. The results show that $\varepsilon$ significantly affects the coherence of the solitons, and its robustness decreases as $|\varepsilon|$ increases. Our research results not only provide a new theoretical basis for understanding soliton dynamics, but also offer important guidance for practical applications, such as optical fiber communication and fluid dynamics.
@@ 1. Introduction
Fluid mechanics plays an important role in engineering and science by verifying theories to predict and analyze complex flow phenomena [1]. This discipline studies fluid motion and static behavior, adheres to Newton’s laws, and establishes parametric scaling models through dimensionless analysis, providing a foundation for interdisciplinary research [2, 3]. Solitons, which preserve their shapes during the propagation [4, 5], hold significant importance in fluid mechanics, particularly in the study of shallow water waves, ocean waves, and internal waves in stratified fluids [3]. Besides, solitons are also extensively studied in nonlinear optics [6, 7, 8], Bose-Einstein condensates [9], plasma physics [10], and molecular biology [11]. Solitons are steady nonlinear waves, while other transient studies focusing on their unsteady forms can be found in Refs. [12, 13]. The focusing nonlinear Schrödinger (NLS) equation governs the propagation of optical solitons in the picosecond regime, balancing group velocity dispersion with self-phase modulation [14]. Experimental observations and theoretical predictions of optical soliton propagation in optical fibers [4, 6, 7] have strongly driven the development of both mathematical and physical models for solitons. The focusing NLS equation is known to admit $N$-soliton solutions, which describe the elastic interactions of $N$ solitons in an ideal optical Kerr medium, where the reflection coefficient has $N$ simple poles according to the inverse scattering transform [15, 16]. Namely, an $N$-soliton solution corresponds to $N$ distinct eigenvalues of the linear spectral problem. As the time evolution $t\to\pm\infty$, an $N$-soliton solution can be viewed as a superposition of $N$ individual solitons, which separate linearly with $t$ and exhibit no interaction force [15, 16]. When the eigenvalues have the same real parts but different imaginary parts, the $N$ interacting solitons have the same velocity, thereby forming a bound state called soliton molecule[14, 17]. These soliton molecules exhibit molecular dynamic properties, and are proposed for increasing the bit rate in multi-level optical communication applications [18, 19].
When the eigenvalues tend to be conjugate or same value, the soliton solution exhibits special dynamical behaviors. In this case, the soliton solution will transform from the ordinary multi-soliton solution to a degenerate soliton solution with more complex behaviors, i.e., multi-pole solution in the context of the inverse scattering transform [16, 20, 21, 22]. Distinct from the usual $N$-soliton interactions, degenerate solitons, referred to as special soliton molecules, can display intense interactions in the near-field region, where the relative distance between interacting solitons increases logarithmically with $|t|$ [21, 22]. Additionally, the initial conditions such as the initial amplitude and phase of the solitons, as well as external disturbances, non-uniform media, or higher-power input, can all trigger degradation phenomena, disrupting the stable equilibrium of the solitons and altering their dynamic behaviors [21, 22].
Degradation phenomena usually occur under specific conditions. For instance, in the coupled Hirota system, when the higher-order dispersion parameter reaches a certain value, the characteristic values of the solitons may tend to be conjugate, thereby triggering degradation [21, 22]. The degradation phenomenon has various impacts on the actual system. In fluid mechanics, degenerate solitons may correspond to complex wave interactions, affecting the transmission and distribution of wave energy [23, 24, 25]. For instance, their strong interactions may lead to energy redistribution, which has potential applications in marine engineering and wave energy utilization. Degenerate soliton solutions serve as the useful models for describing the behavior of multiple chirped pulses with the identical amplitudes and group velocities, especially when they are introduced without any phase difference [23, 24, 25]. Mixed solitons can describe complex interactions among different types of solitons, such as degenerate solitons and bell-shaped solitons, which can reveal new behaviors of solitons in coupled systems [24, 25]. The study of mixed solitons not only helps to investigate complex interaction dynamics, but also provides a more realistic description of multi-component systems, and can incorporate higher-order effects such as third-order dispersion and self-steeping effects, which is crucial for a comprehensive understanding of soliton behavior. In terms of practical applications, the research results of mixed solitons are also of great significance in fields such as nonlinear optics including optical communication systems and optical switch devices, and fluid mechanics including wave behavior and coastal protection measures [24, 25]. In recent years, degenerate soliton and mixed soliton solutions have been investigated in coupled NLS systems, revealing some fascinating dynamical properties [25, 26, 27].
@@
Coupled NLS systems are highly significant in both theory and practice for describing the behavior of multiple modes, frequencies, and polarizations in optical fibers and related structures [28, 29, 30]. These systems enable intensities transfer among additional degrees of freedom and generate a wide variety of vector solutions [28, 29, 30]. To describe the propagation of orthogonally polarized optical waves in an isotropic medium, the coupled NLS system with negative coherent coupling has been studied, which exhibits abundant coherent nonlinear phenomena [31, 32]. Additionally, for higher pulse input powers, it is important to incorporate higher-order effects into the basic NLS systems [14, 33, 34, 35]. In nonlinear optics, the third-order dispersion and self-steepening effects are used to describe the propagation of ultrashort pulses in optical fibers [14, 33, 34, 35]. Consequently, researchers have introduced the coupled Hirota system with higher-order effects which can find applications in nonlinear optics [36, 37, 38]
$\hskip-42.67912pt{\rm i}q_{1,t}+q_{1,xx}+2\left(|q_{1}|^{2}+2|q_{2}|^{2}\right)q_{1}-2q_{1}^{*}q_{2}^{2}+{\rm i}\varepsilon q_{1,xxx}$
$\hskip-28.45274pt+6{\rm i}\varepsilon\left(|q_{1}|^{2}+|q_{2}|^{2}\right)q_{1,x}+6{\rm i}\varepsilon\left(q_{1}q_{2}^{*}-q_{2}q_{1}^{*}\right)q_{2,x}=0,$ (1a)
$\hskip-42.67912pt{\rm i}q_{2,t}+q_{2,xx}+2\left(|q_{2}|^{2}+2|q_{1}|^{2}\right)q_{2}-2q_{2}^{*}q_{1}^{2}+{\rm i}\varepsilon q_{2,xxx}$
$\hskip-28.45274pt+6{\rm i}\varepsilon\left(|q_{1}|^{2}+|q_{2}|^{2}\right)q_{2,x}+6{\rm i}\varepsilon\left(q_{2}q_{1}^{*}-q_{1}q_{2}^{*}\right)q_{1,x}=0,$ (1b)
@@
where $q_{1}=q_{1}(x,t)$ and $q_{2}=q_{2}(x,t)$ represent the slowly varying complex amplitudes of two interacting optical modes, the functions depend on the propagation variable $x$ and the time variable $t$, and the asterisk denotes the complex conjugate. In the context of short-pulse propagation through weakly birefringent media exhibiting Kerr-type nonlinearity, the variable $t$ represents the retarded time. The terms $|q_{1}|^{2}q_{1}$ and $|q_{2}|^{2}q_{2}$ account for self-phase modulation, while the terms $|q_{1}|^{2}q_{2}$ and $|q_{2}|^{2}q_{1}$ describe cross-phase modulation, and the terms $q_{1}^{*}q_{2}^{2}$ and $q_{2}^{*}q_{1}^{2}$ represent coherent coupling effects governing the energy exchange between the two modes of the fiber, and $\varepsilon$ is a higher-order perturbation parameter. When $\varepsilon\neq 0$, System (1) incorporates high-order effects such as third-order dispersion and self-steepening, which facilitate the description of orthogonally polarized optical pulse propagation in an isotropic medium. System (1) can be symplified into the matrix form of
$\hskip 0.0pt{\rm i}{\bf Q}_{t}+{\bf Q}_{xx}+2{\bf QQ}^{\dagger}{\bf Q}+{\rm i}\varepsilon{\bf Q}_{xxx}+6{\rm i}\varepsilon{\bf QQ}^{\dagger}{\bf Q}_{x}={\bf 0},$ (2)
@@
where ${\bf Q}=\begin{pmatrix}q_{1}&q_{2}\\ -q_{2}&q_{1}\end{pmatrix}$ and “ $\dagger$ ” denotes the conjugate transpose, which is also called the matrix Hirota equation. It is noted that System (1) is unchanged through $x\rightarrow-x$ and $\varepsilon\rightarrow-\varepsilon$. Namely, we only need to consider the absolute value of higher-order perturbation parameter $\varepsilon$ and always set $\varepsilon>0$ in this paper.
@@
For System (1), Ref. [36] has explored one- and two-hump solitons, bright and dark rogue waves, and bright and dark breathers with respect to a single spectral parameter using the Darboux transformation (DT) and generalized DT. Ref. [37] has studied the interactions of vector breathers with two spectral parameters, while Ref. [38] has reported the hybrid structures of rogue waves and breathers by the generalized DT. In this paper, we will extend our previous research by investigating degenerate and mixed solitons, giving asymptotic analysis of their interactions, and performing numerical studies on their robustness, all of which are novel contributions not reported in Ref. [37, 38].
In this paper, we focus on the degenerate and mixed solitons for the coupled Hirota system with strong coherent coupling effects. This system can more accurately describe the behavior of solitons with strong coherent coupling effects, especially when higher-order dispersion and self-steepening effects cannot be ignored. Our research results not only can enrich the soliton theory but also provide important guidance for practical applications. For example, in optical fiber communication, understanding the degeneration phenomenon and interaction mechanism of solitons helps to optimize signal transmission schemes, reduce signal distortion and energy loss; in fluid dynamics, a deep understanding of soliton dynamics can provide theoretical basis for the development and utilization of ocean wave energy, and also provide support for predicting and preventing marine disasters. Moreover, the binary Darboux transformation method and asymptotic analysis technique adopted in this paper provide new ideas and methods for solving other complex nonlinear systems, with wide applicability and promotion value.
@@
The aim of this paper is to investigate the asymptotic behaviors and coherent dynamics for the higher-order effects of degenerate and mixed solitons for System (1). The structure of this paper is outlined as follows: In Section 2, we will construct the $N$th-order binary DT and derive the determinant representation of soliton solutions for System (1). Section 3 will focus on deriving three types of degenerate solitons and their asymptotic expressions based on the binary DT. In Section 4, we will analyze the asymptotic expressions of mixed solitons before and after interactions, describing the four interaction mechanisms. In Section 5, we will discuss the interaction properties of mixed solitons based on both asymptotic and graphical analyses. In Section 6, we will analyze a special degenerate vector soliton with significant coherence effects and numerically investigate the relationship between the soliton’s robustness and parameter $\varepsilon$. Section 7 will conclude the paper.
@@ 2. The $N$th-order binary DT
Based on the Ablowitz-Kaup-Newell-Segur inverse scattering formulation [39], System (1) is completely integrable and admit the $\times 4$ 4 Lax pair [36, 37, 38]
$\hskip-56.9055pt\Psi_{x}=U(\lambda;W)\Psi,\qquad\Psi_{t}=V(\lambda;W)\Psi,$ (3)
where $\Psi$ is a 4×14$\times$ 14 × 1 vector complex differentiable eigenfunction of $x$, $t$ and the complex spectral parameter $\lambda$, and $U$ and $V$ have the following forms of
$\hskip-31.2982ptU(\lambda;W)={\rm i}\lambda\Omega+W,$
$\hskip-31.2982ptV(\lambda;W)=4{\rm i}\varepsilon\Omega\lambda^{3}+2\left(2\varepsilon W+{\rm i}\Omega\right)\lambda^{2}+2\left({\rm i}\varepsilon W^{2}\Omega+W+{\rm i}\varepsilon W_{x}\Omega\right)\lambda$
$\hskip-8.5359pt+2\varepsilon W^{3}+{\rm i}W^{2}\Omega-\varepsilon W_{xx}+{\rm i}W_{x}\Omega+\varepsilon\left(W_{x}W-WW_{x}\right),$
@@
$\hskip-31.2982pt\Omega=\begin{pmatrix}-{\bf I}_{2\times 2}&{\bf 0}_{2\times 2}\\ {\bf 0}_{2\times 2}&{\bf I}_{2\times 2}\end{pmatrix},\qquad W=\begin{pmatrix}{\bf 0}_{2\times 2}&{\bf Q}\\ -{\bf Q}^{{\dagger}}&{\bf 0}_{2\times 2}\end{pmatrix},$
${\bf I}_{2\times 2}$ and ${\bf 0}_{2\times 2}$ as the $\times 2$ 2 identity matrix and zero matrix, respectively. From Lax Pair (3), we find that the compatibility condition $U_{t}-V_{x}+UV-VU=0$ is equivalent to System (1).
@@
[…]
$\hskip-28.45274pt\Gamma_{jx}=U(\lambda_{j};W)\Gamma_{j},\qquad\Gamma_{jt}=V(\lambda_{j};W)\Gamma_{j}.$ (5)
@@
[…]
$\hskip-28.45274pt\varUpsilon_{1}[N]=\left(\begin{array}[]{ccccccc}\psi_{11}&-\psi_{12}&\psi_{21}&-\psi_{22}&\cdots&\psi_{N,1}&-\psi_{N,2}\\ \psi_{12}&\psi_{11}&\psi_{22}&\psi_{21}&\cdots&\psi_{N,2}&\psi_{N,1}\end{array}\right)=\left(\mathcal{H}_{1},\mathcal{H}_{2},\cdots,\mathcal{H}_{N}\right),$
$\hskip-28.45274pt\varUpsilon_{2}[N]=\left(\begin{array}[]{ccccccc}\psi_{13}&-\psi_{14}&\psi_{23}&-\psi_{24}&\cdots&\psi_{N,3}&-\psi_{N,4}\\ \psi_{14}&\psi_{13}&\psi_{24}&\psi_{23}&\cdots&\psi_{N,4}&\psi_{N,3}\end{array}\right)=\left(\mathcal{Y}_{1},\mathcal{Y}_{2},\cdots,\mathcal{Y}_{N}\right),$
@@
$\hskip-28.45274pt\varOmega[N]=\left(\begin{array}[]{cccc}\varOmega\left(\Gamma_{1},\Gamma_{1}\right)&\varOmega\left(\Gamma_{1},\Gamma_{2}\right)&\cdots&\varOmega\left(\Gamma_{1},\Gamma_{N}\right)\\ \varOmega\left(\Gamma_{2},\Gamma_{1}\right)&\varOmega\left(\Gamma_{2},\Gamma_{2}\right)&\cdots&\varOmega\left(\Gamma_{2},\Gamma_{N}\right)\\ \vdots&\vdots&\ddots&\vdots\\ \varOmega\left(\Gamma_{N},\Gamma_{1}\right)&\varOmega\left(\Gamma_{N},\Gamma_{2}\right)&\cdots&\varOmega\left(\Gamma_{N},\Gamma_{N}\right)\\ \end{array}\right),$ (13)